By using the ideal gas equation of state, show that change in the internal energy of the same quantity of an ideal gas in adiabatic process is given by
dU= -NRTlnV
1
Expert's answer
2019-08-29T09:42:19-0400

We can write the first law of thermodynamics as

"Q = \u0394U +W (1)"

where Q is the heat supplied to the system ΔU is the internal energy of a closed system, W is the work done by the system on its surroundings


We can write for adiabatic process

"Q =0 (2)"


We put (2) in (1)

"0 = \u0394U +W (3)"

The ideal gas equation of state for isothermal state as

"W= NRT \ln{V} (4)"

Using (3) and (4) we got

"dU= - NRT \ln{V}"